# Gravitational Waves Emission in Quadratic Gravity: longitudinal modes, angular momentum emission, and positivity of the radiated power

### v1

@@ III.2 Total power radiated
$-2\left(q_{n}^{\Psi}\right)^{2}\left(\dot{\Psi}_{B}^{\left(n\right)}\right)^{2}$
$+2\left(\dot{\Psi}_{C}^{\left(n\right)}\right)^{2}-2\left(q_{n}^{\Psi}\right)^{2}\left(\dot{\Psi}_{C}^{\left(n\right)}\right)^{2}$
$+\frac{3}{2}\left(1-\left(q_{n}^{\Psi}\right)^{2}\right)^{2}\left(\dot{\Psi}_{D}^{\left(n\right)}\right)^{2}$
Eqs. (139) and (III.2) can be inserted into Eq. (78). This yields the expression of the total power radiated in terms of the propagation modes:
@@
$-q_{n}^{\Psi}\left(1-\left(q_{n}^{\Psi}\right)^{2}\right)\left[\left\langle\left(\dot{\Psi}_{B}^{\left(n\right)}\right)^{2}\right\rangle+\left\langle\left(\dot{\Psi}_{C}^{\left(n\right)}\right)^{2}\right\rangle\right]\Theta_{n}^{\Psi}$ (141)
$\left.-\frac{3}{4}q_{n}^{\Psi}\left(1-\left(q_{n}^{\Psi}\right)^{2}\right)^{2}\left\langle\left(\dot{\Psi}_{D}^{\left(n\right)}\right)^{2}\right\rangle\Theta_{n}^{\Psi}\right\}.$
@@
The integration with respect to the angular variables in $d\Omega$ in Eq. (141) instruct us to account for the various modes propagating in an arbitrary direction. By inserting Eqs. (117), (118), (119), (120), (121), (122), (123), and (124) into (141), we obtain, after a long calculation, the following:
$P$ $=\frac{G}{5c^{5}}\sum_{n}\left\{\left\langle\dddot{Q}_{ij}^{\left(n\right)}\dddot{Q}_{ij}^{\left(n\right)}\right\rangle\right.$ (142)
@@
$\left.\left.+\frac{1}{2}\left(\frac{m_{\Psi}c}{\omega_{n}}\right)^{4}\left\langle\left(1-\hat{O}_{\Psi}\right)\dddot{K}_{ij}^{\left(n\right)}\left(1-\hat{O}_{\Psi}\right)\dddot{K}_{ij}^{\left(n\right)}\right\rangle\right]\Theta_{n}^{\Psi}\right\},$
@@
We track the contribution of each propagation mode to the various term in Eq. (142) and establish the following association
@@
$q_{n}^{\Psi}\left(\left\langle\left(1-\hat{O}_{\Psi}\right)\dddot{Q}_{ij}^{\left(n\right)}\left(1-\hat{O}_{\Psi}\right)\dddot{Q}_{ij}^{\left(n\right)}\right\rangle\right)$ $\rightarrow\Psi_{+,\times},$
$q_{n}^{\Psi}\left(\frac{m_{\Psi}c}{\omega_{n}}\right)^{2}\left(\left\langle\left(1-\hat{O}_{\Psi}\right)\dddot{Q}_{ij}^{\left(n\right)}\left(1-\hat{O}_{\Psi}\right)\dddot{Q}_{ij}^{\left(n\right)}\right\rangle\right)$ $\rightarrow\Psi_{B,C},$
$q_{n}^{\Psi}\left(\frac{m_{\Psi}c}{\omega_{n}}\right)^{4}\left(\left\langle\left(1-\hat{O}_{\Psi}\right)\dddot{K}_{ij}^{\left(n\right)}\left(1-\hat{O}_{\Psi}\right)\dddot{K}_{ij}^{\left(n\right)}\right\rangle\right)$ $\rightarrow\Psi_{D},$
@@
The sign of $P$ is the concerning point regarding Eq. (149). Its first line is positive definite, since $0\leq q_{n}^{X}\leq 1$ (with $X=\Psi$,$\Phi$). In fact, the modes $\left\{h_{+},\Psi_{+}\right\}$ and $\left\{h_{\times},\Psi_{\times}\right\}$ combine in a pattern of destructive interference. The same is not true for the term in the second line of Eq. (149) with relation to the terms involving the longitudinal modes $\Psi_{B}$, $\Psi_{C}$, and $\Psi_{D}$. The contribution of the latter modes for $P$ will always be negative.
@@ III.3 Binary point-mass system in circular orbit
One of the simplest (and most interesting) physical systems to be studied is the set of two point masses $m_{1}$ and $m_{2}$ in mutual circular orbit with radius $R$ and angular frequency $\omega_{s}$. The trajectories of both particles define a plane of motion, which may be called the $xy$-plane. Accordingly, the non-null mass moments are $M_{11}=-M_{22}$ and $M_{12}$, only [47]. Under the assumption of a non-relativistic motion, it can be shown that the emitted gravitational wave is monochromatic of frequency $\omega=2\omega_{s}$; it is also true that [40]:
@@
$P$ $=$ $P_{\text{GR}}\left[1-\sqrt{1-\left(\frac{m_{\Psi}c}{2\omega_{s}}\right)^{2}}\left(1+\left(\frac{m_{\Psi}c}{2\omega_{s}}\right)^{2}\right.\right.$ (152)
$\left.\left.+\frac{1}{2}\left(\frac{m_{\Psi}c}{2\omega_{s}}\right)^{4}\right)\right],$
@@
where $E_{\text{orbit}}$ is the Newtonian orbital energy [^7] The non-relativistic binary black hole systems are described by Newtonian dynamics in quadratic gravity [49].
@@
$\dot{\omega}_{s}$ $=$ $\omega_{\text{GR}}\left[1-\sqrt{1-\left(\frac{m_{\Psi}c}{2\omega_{s}}\right)^{2}}\left(1+\left(\frac{m_{\Psi}c}{2\omega_{s}}\right)^{2}\right.\right.$ (157)
$\left.\left.+\frac{1}{2}\left(\frac{m_{\Psi}c}{2\omega_{s}}\right)^{4}\right).\right],$
@@
The main distinction between the curves displayed in Fig. 2 is the following. The angular frequency as a function of time in GR is the black curve; it diverges in the finite time $\frac{t}{t_{\text{ref}}}=\frac{3}{8}$. The blue curve gives $\omega_{s}\left(t\right)$ for the complete spin-2 mode; it approaches the constant value $\omega_{s}^{\max}\simeq 1.145\omega_{0}$ as $t\rightarrow\infty$. The finite value of the angular frequency for the complete case occurs in the limit of the square bracket of Eq. (157) approaching zero, i.e. for the power $P=0$. Therefore, in this particular case, the total power radiated will never assume negative values.
@@
$h_{+}\left(\mathbf{\hat{z}}\right)$ $=\frac{\kappa}{4\pi r}\ddot{M}_{11}$ (159)
$=A_{\max}\cos\left[2\omega_{s}^{\max}\left(t-\frac{r}{c}\right)\right],$
$h_{\times}\left(\mathbf{\hat{z}}\right)$ $=\frac{\kappa}{4\pi r}\ddot{M}_{12}$ (160)
$=A_{\max}\sin\left[2\omega_{s}^{\max}\left(t-\frac{r}{c}\right)\right],$
@@
$\Psi_{+}\left(\mathbf{\hat{z}}\right)$ $=\frac{\kappa}{4\pi r}\left[1-\hat{O}_{\Psi}\right]\ddot{M}_{11}$ (161)
$=-A_{\max}\cos\left[2\omega_{s}^{\max}\left(t-\left(\frac{r}{c}\right)q_{\max}^{\Psi}\right)\right],$
$\Psi_{\times}\left(\mathbf{\hat{z}}\right)$ $=\frac{\kappa}{4\pi r}\left[1-\hat{O}_{\Psi}\right]\ddot{M}_{12}$ (162)
$=-A_{\max}\sin\left[2\omega_{s}^{\max}\left(t-\left(\frac{r}{c}\right)q_{\max}^{\Psi}\right)\right],$
@@
The results in Eqs. (159), (160), (161), and (162) were explicitly obtained in Ref. [40].
@@
The previous analysis clearly shows a physical inconsistency in the modified gravitational model even with $P$ being always positive valued. The next section makes it clear that this caveat is related to the presence of longitudinal modes in the quadrupole approximation.
@@ III.4 Transverse-traceless spin22-2- 2 massive modes
The problem revealed in the previous sections is incepted by the term $C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta}$ in the Lagrangian of the quadratic gravity. In fact, it is long known in the literature [17] that such a term presents an Ostrogradsky instability leading to non-positive definite energy. In the context of gravitational waves, this translates to negative-valued radiated power, a clear physical inconsistency.
One way to remedy this issue is to argue that the massive spin-2 mode will manifest itself exclusively in the damping regime thus avoiding pathological propagating modes. This is equivalent to state that $m_{\Psi}$ very large so that $\frac{c}{\omega}>m_{\Psi}^{-1}$ in the classical length scales. In this way, the term $C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta}$ will be important in the quantum domain while being completely negligible in the classical context of gravitational waves. [^8] In the quantum realm, Ostrogradsky instability translates to the existence of ghost-like fields. Even though ghost fields are problematic [31], there are approaches in the literature dedicated to dealing consistently with this type of fields [33, 34, 35, 36, 50, 51].
Another way to deal with this problem is to impose that the transverse-traceless massive spin-2 modes are the only modes present in the theory. In other words, there should be
@@
$-q_{n}^{\Psi}\Theta_{n}^{\Psi}\left\langle\dot{\Psi}_{ab}^{TT}\dot{\Psi}_{ab}^{TT}\right\rangle_{\left(n\right)}+\left.6q_{n}^{\Phi}\Theta_{n}^{\Phi}\left\langle\dot{\Phi}^{2}\right\rangle_{\left(n\right)}\right],$
@@
$\frac{dJ^{i}}{dt}$ $=\frac{c^{3}r^{2}}{32\pi G}\sum_{n}\epsilon^{ikl}\int d\Omega\left\{-\left\langle\dot{h}_{ab}^{TT}x_{k}\partial_{l}h_{ab}^{TT}\right\rangle_{\left(n\right)}\right.$ (169)
@@
$\left.-2q_{n}^{\Psi}\Theta_{n}^{\Psi}\left\langle\Psi_{ak}^{TT}\dot{\Psi}_{al}^{TT}\right\rangle_{\left(n\right)}-6q_{n}^{\Phi}\Theta_{n}^{\Phi}\left\langle\dot{\Phi}x_{k}\partial_{l}\Phi\right\rangle_{\left(n\right)}\right\}.$
@@
The description of a binary system of point-like masses in circular orbits under the prescription (167) was studied by [40] in detail. In that reference, the physical consistency of the system was shown in the context where the gravitational waves carry energy away, only. In the next section, we study another example where the gravitational waves emission carries away both energy and angular momentum.
@@ IV Rotating ellipsoid with free precession
This section is devoted to the study of an astrophysical object that may be described as an ellipsoid rotating with free precession. In most astronomical objects, the rotation axis typically does not align with a principal axis. As a result, the motion of the rigid body involves a combination of rotation around a principal axis and precession. As is well explored in [47], this precession motion introduces qualitatively new characteristics into the emitted gravitational waves. As we will see, these waves reduce both the energy and angular momentum of the body.
@@
Consider an axisymmetric body whose longitudinal axis $x_{3}^{\prime}$ forms an angle $\alpha$ with the axis of angular momentum $x_{3}$. This angle $\alpha$ is referred to as the wobble angle and the gravitational wave emission associated with this motion is known as wobble radiation [47]. When the rotation axis does not align with the principal axis of the body, the angular momentum rotates around the principal axis, leading to precession. According to [52], by solving the force-free Euler equations, it is evident that free precession occurs as the principal axis rotates around the angular momentum axis with a frequency of
@@
where the angular momentum is $\vec{J}=\left(0,0,J\right)$. In Section III.1, we derived the propagation modes for gravitational waves in a generic direction, considering all polarizations of the fields $h_{\mu\nu}$, $\Psi_{\mu\nu}$, and $\Phi$. Additionally, in Section III.4, we argued that in order to properly address the pathological propagation modes of the massive spin-$2$ field, it is necessary to impose that only the transverse-traceless massive spin-$2$ modes are present in the theory. With these considerations in mind, and utilizing $\phi=0$, (i) equations (117) and (118) for the massless spin-$2$ modes; (ii) equations (119) and (120) for the massive spin-$2$ modes, and (iii) (124) for the massive spin-00 mode we obtain
@@
    $A_{s,1}=\Phi_{0}^{\prime}\sin\theta\cos\theta\sin 2\alpha,$ (188)
@@
The resulting gravitational waves for all fields are a combination of two frequency spectra: $\Omega$ and $2\Omega$. The presence of these two spectra provides the following interpretation for the damping and oscillatory modes of the massive fields: Initially, when $\Omega<m_{X}c$, both waves (with frequencies $\omega_{gw}=\Omega$ and $\omega_{gw}=2\Omega$) are in damping modes. As $\Omega$ increases to $\Omega>\frac{m_{X}c}{2}$, the wave with frequency $\omega_{gw}=2\Omega$ transitions to the oscillatory mode, while the wave with $\omega_{gw}=\Omega$ remains in the damping mode. Finally,when $\Omega>m_{X}c$, the wave with $\omega_{gw}=\Omega$ also enters the oscillatory mode, completing the waveform for the corresponding massive mode. Note that in the full oscillatory regime, the solutions $h_{+,\times}$ and $\Psi_{+,\times}$ are waves with the same amplitude and frequency but with different wave numbers. This allows for an interpretation of destructive interference effects, as discussed in [40]. Furthermore, in the limit of $m_{\Psi}\rightarrow 0$, this destructive interference becomes complete, resulting in the absence of tensorial mode emission.
Using the equations above, we can calculate the loss of energy and angular momentum through Eqs. (III.4) and (169). By substituting the propagation modes into (III.4), we obtain the energy loss as
$\frac{dE}{dt}$ $=A_{E}\left\{16\sin^{2}\alpha\left[1-\sqrt{1-\left(\frac{m_{\Psi}c}{2\Omega}\right)^{2}}\right.\right.$ (189)
@@
$\left.\left.+\frac{1}{18}\sqrt{1-\left(\frac{m_{\Phi}c}{\Omega}\right)^{2}}\right]\right\},$
@@
$\frac{dJ}{dt}$ $=A_{J}\left\{16\sin^{2}\alpha\left[1-\sqrt{1-\left(\frac{m_{\Psi}c}{2\Omega}\right)^{2}}\right.\right.$ (191)
@@
$\left.\left.+\frac{1}{18}\sqrt{1-\left(\frac{m_{\Phi}c}{\Omega}\right)^{2}}\right]\right\},$
@@
$\left.\left.+\frac{1}{18}\sqrt{1-\left(\frac{m_{\Phi}c}{\Omega}\right)^{2}}\right]\right\},$
@@
$\dot{\alpha}$ $=A_{\alpha}\left\{16\sin^{2}\alpha\left[1-\sqrt{1-\left(\frac{m_{\Psi}c}{2\dot{\beta}}\right)^{2}}\right.\right.$ (195)
@@
$\left.\left.+\frac{1}{18}\sqrt{1-\left(\frac{m_{\Phi}c}{\dot{\beta}}\right)^{2}}\right]\right\},$
@@
In order to examine the coupled system given by Eqs. (193) and (195) in greater detail, we introduce the dimensionless function is given by

### v2

@@ III.2 Total power radiated
$-2\left(q_{n}^{\Psi}\right)^{2}\left(\dot{\Psi}_{B}^{\left(n\right)}\right)^{2}-2\left(q_{n}^{\Psi}\right)^{2}\left(\dot{\Psi}_{C}^{\left(n\right)}\right)^{2}$
$+2\left(\dot{\Psi}_{C}^{\left(n\right)}\right)^{2}+\frac{3}{2}\left(1-\left(q_{n}^{\Psi}\right)^{2}\right)^{2}\left(\dot{\Psi}_{D}^{\left(n\right)}\right)^{2}.$ (136)
Eqs. (135) and (136) can be inserted into Eq. (75). This yields the expression of the total power radiated in terms of the propagation modes:
@@
$-q_{n}^{\Psi}\left(1-\left(q_{n}^{\Psi}\right)^{2}\right)\left[\left\langle\left(\dot{\Psi}_{B}^{\left(n\right)}\right)^{2}\right\rangle+\left\langle\left(\dot{\Psi}_{C}^{\left(n\right)}\right)^{2}\right\rangle\right]\Theta_{n}^{\Psi}$
$\left.-\frac{3}{4}q_{n}^{\Psi}\left(1-\left(q_{n}^{\Psi}\right)^{2}\right)^{2}\left\langle\left(\dot{\Psi}_{D}^{\left(n\right)}\right)^{2}\right\rangle\Theta_{n}^{\Psi}\right\}.$ (137)
@@
The integration with respect to the angular variables in $d\Omega$ in Eq. (137) instructs us to account for the various modes propagating in an arbitrary direction. By inserting Eqs. (111), (112), (113), (114), (115), (116), (117), and (118) into (137), we obtain, after a long calculation, the following:
$P=\frac{G}{5c^{5}}\sum_{n}\left\{\left\langle\dddot{Q}_{ij}^{\left(n\right)}\dddot{Q}_{ij}^{\left(n\right)}\right\rangle\right.$
@@
$\left.\left.+\frac{1}{2}\left(\frac{m_{\Psi}c}{\omega_{n}}\right)^{4}\left\langle\left(1-\hat{O}_{\Psi}\right)\dddot{K}_{ij}^{\left(n\right)}\left(1-\hat{O}_{\Psi}\right)\dddot{K}_{ij}^{\left(n\right)}\right\rangle\right]\Theta_{n}^{\Psi}\right\},$ (138)
@@
We track the contribution of each propagation mode to the various terms in Eq. (138) and establish the following association
@@
$q_{n}^{\Psi}\left\langle\left(1-\hat{O}_{\Psi}\right)\dddot{Q}_{ij}^{\left(n\right)}\left(1-\hat{O}_{\Psi}\right)\dddot{Q}_{ij}^{\left(n\right)}\right\rangle\rightarrow\Psi_{+,\times},$
$q_{n}^{\Psi}\left(\frac{m_{\Psi}c}{\omega_{n}}\right)^{2}\left\langle\left(1-\hat{O}_{\Psi}\right)\dddot{Q}_{ij}^{\left(n\right)}\left(1-\hat{O}_{\Psi}\right)\dddot{Q}_{ij}^{\left(n\right)}\right\rangle\rightarrow\Psi_{B,C},$
$q_{n}^{\Psi}\left(\frac{m_{\Psi}c}{\omega_{n}}\right)^{4}\left\langle\left(1-\hat{O}_{\Psi}\right)\dddot{K}_{ij}^{\left(n\right)}\left(1-\hat{O}_{\Psi}\right)\dddot{K}_{ij}^{\left(n\right)}\right\rangle\rightarrow\Psi_{D},$
@@
The sign of $P$ is the concerning point regarding Eq. (145). Its first two lines are positive definite, since $0\leq q_{n}^{X}\leq 1$ (with $X=\Psi$,$\Phi$). In fact, the modes $\left\{h_{+},\Psi_{+}\right\}$ and $\left\{h_{\times},\Psi_{\times}\right\}$ combine in a pattern of destructive interference. The same is not true for the terms in the last two lines of Eq. (145) with relation to the terms involving the longitudinal modes $\Psi_{B}$, $\Psi_{C}$, and $\Psi_{D}$. The contribution of the latter modes for $P$ will always be negative.
@@ III.3 Binary point-mass system in circular orbit
One of the simplest (and most interesting) physical systems to be studied is the set of two point masses $m_{1}$ and $m_{2}$ in mutual circular orbit with radius $R$ and angular frequency $\omega_{s}$. The trajectories of both particles define a plane of motion, which may be called the $xy$-plane. Accordingly, the non-null mass moments are $M_{11}=-M_{22}$ and $M_{12}$, only Maggiore (2007). Under the assumption of a non-relativistic motion, it can be shown that the emitted gravitational wave is monochromatic of frequency $\omega=2\omega_{s}$; it is also true that Alves et al. (2023):
@@
$P$ $=P_{\text{GR}}\left[1-\sqrt{1-\left(\frac{m_{\Psi}c}{2\omega_{s}}\right)^{2}}\left(1+\left(\frac{m_{\Psi}c}{2\omega_{s}}\right)^{2}\right.\right.$
$\qquad\quad\left.\left.+\frac{1}{2}\left(\frac{m_{\Psi}c}{2\omega_{s}}\right)^{4}\right)\right],$ (148)
@@
where $E_{\text{orbit}}$ is the Newtonian orbital energy [^7] The non-relativistic binary black hole systems are described by Newtonian dynamics in quadratic gravity Lü et al. (2015).
@@
$\dot{\omega}_{s}$ $=\omega_{\text{GR}}\left\{1-\sqrt{1-\left(\frac{m_{\Psi}c}{2\omega_{s}}\right)^{2}}\left[1+\left(\frac{m_{\Psi}c}{2\omega_{s}}\right)^{2}\right.\right.$
$\qquad\quad\left.\left.+\frac{1}{2}\left(\frac{m_{\Psi}c}{2\omega_{s}}\right)^{4}\right]\right\},$ (153)
@@
The main distinction between the curves displayed in Fig. 2 is the following: the angular frequency as a function of time in GR is the black curve; it diverges in the finite time $\frac{t}{t_{\text{ref}}}=\frac{3}{8}$. The blue curve gives $\omega_{s}\left(t\right)$ for the complete spin-2 mode; it approaches the constant value $\omega_{s}^{\max}\simeq 1.145\omega_{0}$ as $t\rightarrow\infty$. The finite value of the angular frequency for the complete case occurs in the limit of the square bracket of Eq. (153) approaching zero, i.e. for the power $P=0$. Therefore, in this particular case, the total power radiated will never assume negative values.
@@
$h_{+}\left(\mathbf{\hat{z}}\right)$ $=\frac{\kappa}{4\pi r}\ddot{M}_{11}$
$=A_{\max}\cos\left[2\omega_{s}^{\max}\left(t-\frac{r}{c}\right)\right],$ (155)
$h_{\times}\left(\mathbf{\hat{z}}\right)$ $=\frac{\kappa}{4\pi r}\ddot{M}_{12}$
$=A_{\max}\sin\left[2\omega_{s}^{\max}\left(t-\frac{r}{c}\right)\right],$ (156)
@@
$\Psi_{+}\left(\mathbf{\hat{z}}\right)$ $=\frac{\kappa}{4\pi r}\left[1-\hat{O}_{\Psi}\right]\ddot{M}_{11}$
$=-A_{\max}\cos\left[2\omega_{s}^{\max}\left(t-\left(\frac{r}{c}\right)q_{\max}^{\Psi}\right)\right],$ (157)
$\Psi_{\times}\left(\mathbf{\hat{z}}\right)$ $=\frac{\kappa}{4\pi r}\left[1-\hat{O}_{\Psi}\right]\ddot{M}_{12}$
$=-A_{\max}\sin\left[2\omega_{s}^{\max}\left(t-\left(\frac{r}{c}\right)q_{\max}^{\Psi}\right)\right],$ (158)
@@
The results in Eqs. (155), (156), (157), and (158) were explicitly obtained in Ref. Alves et al. (2023).
@@
The previous analysis clearly shows a physical inconsistency in the modified gravitational model even with $P$ being always positive valued. This inconsistency also arises in the calculation of the radiated angular momentum given by Eq. (83). Notably, the sign difference between the terms in Eqs. (79) and (80) highlights the presence of Ostrogradsky instability in $j_{\Psi}^{i\left(n\right)}$. Furthermore, the term $\dot{\Psi}_{\left(n\right)}^{\alpha\beta}x_{k}\partial_{l}\Psi_{\alpha\beta}^{\left(n\right)}$ in Eq. (80) contains both transverse and longitudinal modes. Consequently, even when considering the term $\dot{h}_{ab}^{\left(n\right)}x_{k}\partial_{l}h_{ab}^{\left(n\right)}$ in Eq. (79), it is not possible to combine $h_{ab}^{\left(n\right)}$ and $\Psi_{\alpha\beta}^{\left(n\right)}$ in a way that ensures the positivity of Eq. (83).
The next section makes it clear that the issue described previously is related to the presence of longitudinal modes in the quadrupole approximation.
@@ III.4 Transverse-traceless spin22-2- 2 massive modes
The problem revealed in the previous sections is incepted by the term $C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta}$ in the Lagrangian of the quadratic gravity. In fact, it is long known in the literature Stelle (1977) that such a term presents an Ostrogradsky instability leading to non-positive definite energy. In the context of gravitational waves, this translates to negative-valued radiated power, a clear physical inconsistency.
One way to remedy this issue is to argue that the massive spin-2 mode will manifest itself exclusively in the damping regime thus avoiding pathological propagating modes. This is equivalent to state that $m_{\Psi}$ very large so that $\frac{c}{\omega}>m_{\Psi}^{-1}$ in the classical length scales. In this way, the term $C_{\mu\nu\alpha\beta}C^{\mu\nu\alpha\beta}$ will be important in the quantum domain while being completely negligible in the classical context of gravitational waves. [^8] In the quantum realm, Ostrogradsky instability translates to the existence of ghost-like fields. Even though ghost fields are problematic Sbisà (2015), there are approaches in the literature dedicated to dealing consistently with this type of field Modesto and Shapiro (2016); Donoghue and Menezes (2019); Salvio (2019, 2021); ”Anselmi (2017, 2018). Another way to deal with this problem is to impose that the transverse-traceless massive spin-2 modes are the only modes present in the theory. In other words, there should be
@@
$-q_{n}^{\Psi}\Theta_{n}^{\Psi}\left\langle\dot{\Psi}_{ab}^{TT}\dot{\Psi}_{ab}^{TT}\right\rangle_{\left(n\right)}+\left.6q_{n}^{\Phi}\Theta_{n}^{\Phi}\left\langle\dot{\Phi}^{2}\right\rangle_{\left(n\right)}\right],$ (164)
@@
$\frac{dJ^{i}}{dt}$ $=\frac{c^{3}r^{2}}{32\pi G}\sum_{n}\epsilon^{ikl}\int d\Omega\left\{-\left\langle\dot{h}_{ab}^{TT}x_{k}\partial_{l}h_{ab}^{TT}\right\rangle_{\left(n\right)}\right.$
@@
$\left.-2q_{n}^{\Psi}\Theta_{n}^{\Psi}\left\langle\Psi_{ak}^{TT}\dot{\Psi}_{al}^{TT}\right\rangle_{\left(n\right)}-6q_{n}^{\Phi}\Theta_{n}^{\Phi}\left\langle\dot{\Phi}x_{k}\partial_{l}\Phi\right\rangle_{\left(n\right)}\right\}.$ (165)
@@
The description of a binary system of point-like masses in circular orbits under the prescription (163) was studied by the authors of Ref. Alves et al. (2023) in detail. In that reference, the physical consistency of the system was shown in the context where the gravitational waves carry energy away, only. In the next section, we study another example where the gravitational waves emission carries away both energy and angular momentum.
@@ IV Rotating ellipsoid with free precession
This section is devoted to the study of an astrophysical object that may be described as an ellipsoid rotating with free precession. In most astronomical objects, the rotation axis typically does not align with a principal axis. As a result, the motion of the rigid body involves a combination of rotation around a principal axis and precession. As is well explored in Maggiore (2007), this precession motion introduces qualitatively new characteristics into the emitted gravitational waves. As we will see, these waves reduce both the energy and angular momentum of the body.
@@
Consider an axisymmetric body whose longitudinal axis $x_{3}^{\prime}$ forms an angle $\alpha$ with the axis of angular momentum $x_{3}$. This angle $\alpha$ is referred to as the wobble angle and the gravitational wave emission associated with this motion is known as wobble radiation Maggiore (2007). When the rotation axis does not align with the principal axis of the body, the angular momentum rotates around the principal axis, leading to precession. According to Thornton and Marion (2004), by solving the force-free Euler equations, it is evident that free precession occurs as the principal axis rotates around the angular momentum axis with a frequency of
@@
where the angular momentum is $\vec{J}=\left(0,0,J\right)$. In Section III.1, we derived the propagation modes for gravitational waves in a generic direction, considering all polarizations of the fields $h_{\mu\nu}$, $\Psi_{\mu\nu}$, and $\Phi$. Additionally, in Section III.4, we argued that in order to properly address the pathological propagation modes of the massive spin-$2$ field, it is necessary to impose that only the transverse-traceless massive spin-$2$ modes are present in the theory. With these considerations in mind, and utilizing $\phi=0$, (i) equations (111) and (112) for the massless spin-$2$ modes; (ii) equations (113) and (114) for the massive spin-$2$ modes, and (iii) (118) for the massive spin-00 mode, we obtain
@@
    $A_{s,1}=\Phi_{0}^{\prime}\sin\theta\cos\theta\sin 2\alpha,$
@@
The resulting gravitational waves for all fields are a combination of two frequency spectra: $\Omega$ and $2\Omega$. The presence of these two spectra provides the following interpretation for the damping and oscillatory modes of the massive fields: Initially, when $\Omega<m_{X}c$, both waves (with frequencies $\omega_{gw}=\Omega$ and $\omega_{gw}=2\Omega$) are in damping modes. As $\Omega$ increases to $\Omega>\frac{m_{X}c}{2}$, the wave with frequency $\omega_{gw}=2\Omega$ transitions to the oscillatory mode, while the wave with $\omega_{gw}=\Omega$ remains in the damping mode. Finally,when $\Omega>m_{X}c$, the wave with $\omega_{gw}=\Omega$ also enters the oscillatory mode, completing the waveform for the corresponding massive mode. Note that in the full oscillatory regime, the solutions $h_{+,\times}$ and $\Psi_{+,\times}$ are waves with the same amplitude and frequency but with different wave numbers. This allows for an interpretation of destructive interference effects, as discussed in Alves et al. (2023). Furthermore, in the limit of $m_{\Psi}\rightarrow 0$, this destructive interference becomes complete, resulting in the absence of tensorial mode emission.
Using the equations above, we can calculate the loss of energy and angular momentum through Eqs. (164) and (165). By substituting the propagation modes into (164), we obtain the energy loss as
$\frac{dE}{dt}$ $=A_{E}\left\{16\sin^{2}\alpha\left[1-\sqrt{1-\left(\frac{m_{\Psi}c}{2\Omega}\right)^{2}}\right.\right.$
@@
$\left.\left.+\frac{1}{18}\sqrt{1-\left(\frac{m_{\Phi}c}{\Omega}\right)^{2}}\right]\right\},$ (184)
@@
$\frac{dJ}{dt}$ $=A_{J}\left\{16\sin^{2}\alpha\left[1-\sqrt{1-\left(\frac{m_{\Psi}c}{2\Omega}\right)^{2}}\right.\right.$
@@
$\left.\left.+\frac{1}{18}\sqrt{1-\left(\frac{m_{\Phi}c}{\Omega}\right)^{2}}\right]\right\},$ (186)
@@
$\left.\left.+\frac{1}{18}\sqrt{1-\left(\frac{m_{\Phi}c}{\Omega}\right)^{2}}\right]\right\},$ (188)
@@
$\dot{\alpha}$ $=A_{\alpha}\left\{16\sin^{2}\alpha\left[1-\sqrt{1-\left(\frac{m_{\Psi}c}{2\dot{\beta}}\right)^{2}}\right.\right.$
@@
$\left.\left.+\frac{1}{18}\sqrt{1-\left(\frac{m_{\Phi}c}{\dot{\beta}}\right)^{2}}\right]\right\},$ (190)
@@
In order to examine the coupled system given by Eqs. (188) and (190) in greater detail, we introduce the dimensionless function as given by
@@
$\left.\left.+\frac{1}{18}\sqrt{1-\left(\frac{m_{\Phi}c}{\omega\dot{\beta}_{0}}\right)^{2}}\right]\right\},$ (194)